# Multimodal Situational Safety

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) are rapidly evolving, demonstrating impressive capabilities as multimodal assistants that interact with both humans and their environments. However, this increased sophistication introduces significant safety concerns. In this paper, we present the first evaluation and analysis of a novel safety challenge termed Multimodal Situational Safety, which explores how safety considerations vary based on the specific situation in which the user or agent is engaged. We argue that for an MLLM to respond safely—whether through language or action—it often needs to assess the safety implications of a language query within its corresponding visual context. To evaluate this capability, we develop the Multimodal Situational Safety benchmark (MSSBench) to assess the situational safety performance of current MLLMs. The dataset comprises 1,820 language query-image pairs, half of which the image context is safe, and the other half is unsafe. We also develop an evaluation framework that analyzes key safety aspects, including explicit safety reasoning, visual understanding, and, crucially, situational safety reasoning. Our findings reveal that current MLLMs struggle with this nuanced safety problem in the instruction-following setting and struggle to tackle these situational safety challenges all at once, highlighting a key area for future research. Furthermore, we develop multi-agent pipelines to coordinately solve safety challenges, which shows consistent improvement in safety over the original MLLM response. Code and data: mssbench.github.io.
@@ 1 Introduction
To comprehensively evaluate the current MLLM’s situational safety performance, we introduce a new Multimodal Situational Safety benchmark (MSSBench) with 1820 language-image pairs. To assess unbalanced model behaviors, in half of the data, the image is a safe situation for answering the query, and in the other half, the image context is unsafe. Our benchmark considers two multimodal assistant scenarios: multimodal chat agents that respond to the user with their knowledge and multimodal embodied agents that take action to complete household tasks. For the chat scenario, we leverage LLMs to generate candidate activities as user intents and envision an unsafe situation for these activities. Then, the examples will go through two filtering processes: LLM automatic verification and human verification performed by domain experts to ensure data quality. Finally, we prompt the LLMs to generate user queries with the intent to perform these activities. For embodied scenarios, we first manually create potentially unsafe household tasks, and define safe and unsafe situations. Then, we collect safe and unsafe visual contexts from the embodied AI simulators.
We evaluate popular open-sourced and proprietary MLLMs on the MSSBench. The results show that current MLLMs struggle with recognizing unsafe situations when answering user queries. Then, we create different benchmark variants to analyze key safety aspects of MLLMs, including explicit safety reasoning, visual understanding, and situational safety reasoning. Our main findings include: (1) Explicit safety reasoning can improve the average situational safety performance of MLLMs, but will also introduce over-sensitivity in safe situations. (2) All MLLMs perform poorly in embodied scenarios due to the lack of precise visual understanding and situation safety judgment abilities. (3) Open-source MLLMs ignore safety clues in the image with a much higher frequency than proprietary models. (4) Under settings with more subtasks, the safety performance of MLLMs decreases due to task complexity.
@@ 2 Related Work
The second one is embodied decision-making and task planning (Shridhar et al., 2020; Szot et al., 2024), which requires the models to plan and execute actions with visual input from an indoor environment to complete a given household task (Driess et al., 2023; Yang et al., 2024; Li et al., 2024b; Wang et al., 2024a). This enables the MLLMs to control a robot and make it an embodied assistant. However, the improved abilities of current MLLMs on these tasks and new applications introduce new safety problems, and the safety of MLLMs under multimodal assistant scenarios has not been thoroughly studied.
@@
The generative abilities of LLMs and MLLMs carry the risk of being misused to generate harmful content. Recently, lots of efforts have been put into red-teaming MLLMs (Liu et al., 2023c; Gong et al., 2023; Shayegani et al., 2023; Qi et al., 2024; Luo et al., 2024). However, most of the current benchmarks study the scenarios where the language itself is clearly unsafe and leverage image modality as an attack to trick the MLLMs into answering unsafe queries. Liu et al. (2023c) find that using query-relevant images can attack the MLLMs to answer malicious queries. Gong et al. (2023) propose to embed malicious queries into images and leverage the OCR abilities of MLLMs to induce them to generate harmful responses. Moreover, optimized adversarial images are also used to jailbreaking MLLMs (Shayegani et al., 2023). Besides these, there were also concurrent efforts studying the over-sensitivity of MLLMs (Li et al., 2024c). Different from existing works, we first propose a new safety problem for MLLMs in multimodal assistant applications – multimodal situational safety. Based on this, we collect a benchmark containing chat and embodied scenarios to evaluate the MLLMs’ safety awareness in unsafe scenarios and over-sensitivity in safe scenarios. We also investigate in-depth how far we can leverage MLLMs’ capabilities to improve safety performance.
@@ 3.1 Dataset Overview
I. Physical Harm 628 34.5%
            $\bullet$ Self-harm 320 17.6%
            $\bullet$ Self-harm (Embodied Task) 120 6.5%
            $\bullet$ Other-harm 188 10.3%
II. Property Damage 736 40.4%
            $\bullet$ Public property damage 120 6.6%
            $\bullet$ Personal property damage 116 6.4%
            $\bullet$ Personal property damage (Embodied Task) 500 27.5%
III. Offensive Behavior 268 14.7%
            $\bullet$ Cultural belief violations 28 1.5%
            $\bullet$ Disruptive behaviors 148 7.9%
            $\bullet$ Religious belief infringements 92 5.1%
IV. Illegal Activities 188 10.4%
            $\bullet$ Human-restricting activities 76 4.2%
            $\bullet$ Property-restricting activities 88 4.8%
            $\bullet$ Organism-restricting activities 24 1.3%
@@ 3.3 Embodied Data Collection
We mainly consider three task categories: place an {object in hand} on a {receptacle} (Place), toggle a {receptacle} (Toggle), and drop an {object in hand} (Drop). For each category, we can define different safe and unsafe tasks by changing the objects or receptacles in the placeholder. The environment state that determines whether the task is safe or unsafe in the Drop tasks is the object in the robot’s hand. In the Place task, the environment state is the combination of the object in the robot’s hand and the receptacle. In the Toggle task, the environment state is the objects in the receptacle to be toggled. In total, we define 31 safe tasks and 31 unsafe tasks. Then, we create 5 instruction templates for each task. In total, we have $5\times(31+31)=310$ embodied instructions.
@@
After we determine the {object}, {receptacle} in the task, we run a ``$Pick\_\{object\}and\_Place\{receptacle\}"$ task defined in  Shridhar et al. (2020) with the determined {object} and {receptacle}. For the Place task and the Drop task, we randomly collect two egocentric images after the agent picks up the object and before the agent places the object. For the Toggle task, we collect an egocentric image right after the agent places the object on the receptacle from two different episodes. Therefore, we have 620 samples in total. One data example is shown in Fig. 3 (right).
@@ 3.4 Data Statistics
The Multimodal Situational Safety benchmark consists of a substantial collection of 1820 Image-Query pairs, encompassing two subsets: the embodied assistant subset, which contains 620 pairs sourced from household scenarios, and the chat assistant subset, comprising a larger set of 1200 pairs designed for broader situational QA scenarios. Our dataset is a balance dataset, with half of the data containing safe situations and half containing unsafe situations. The statistical details of the data in the MSSBench are presented in Table. 1.
@@ 4.1 Setup
MLLMs. The MLLMs we benchmark include both open-source models and proprietary models accessible only via API. The open-source MLLMs are: (i) LLaVA-1.6 (Liu et al., 2023b), (ii) MiniGPT-v2 (Chen et al., 2023), (iii) Qwen-VL (Bai et al., 2023), (iv) DeepSeek (Lu et al., 2024) and (v) mPLUG-Owl2 (Ye et al., 2024). We implemented these models with their 7B version and using their default settings. For the proprietary models, we evaluated Claude 3.5 Sonnet, GPT-4o (OpenAI, 2023b), and Gemini Pro-1.5 (Reid et al., 2024).
Evaluation. For the instruction following setting, we use GPT-4o (OpenAI, 2023a) to categorize the response generated by MLLMs into safe and unsafe categories. The categories description is introduced in Tables. 4 and 5 in Sec. A.3. Recent studies, including Hsu et al. (2023); Hackl et al. (2023); Wang et al. (2024b) have underscored GPT-4’s effectiveness and reliability in evaluative roles, including safety classification. After categorization, we use accuracy to evaluate MLLM’s safety performance.
@@
MiniGPT-V2 97.6 2.4 50.0 98.8 0.0 49.4 49.8
Qwen-VL 98.0 3.1 50.6 99.0 0.0 49.5 50.2
mPLUG-Owl2 98.7 2.9 50.8 100 0.3 50.0 50.5
Llava 1.6 99.7 2.5 51.1 100 0.6 50.3 50.7
DeepSeek 98.6 6.7 52.7 99.7 0.0 49.9 51.7
@@ 4.2 Main Results
To begin with, we assess the performance of 8 leading multimodal large language models (MLLMs) on our MSS benchmark, the results are shown in Table. 2. To mimic the chat assistant scenario, we inform the MLLM that the image is its first-person view and the query is from a user staying with it, see the ‘Common Prompt’ in Fig. 4. The full prompt can be found in Sec. A.6. First, a common trend among all the MLLMs is that they tend to comply with and answer users’ queries in both safe and unsafe scenarios. This leads to a high safety accuracy when the situation is safe for the user’s intent and a low accuracy when the situation is unsafe. Second, comparing open-source models and proprietary models, we find that proprietary models perform better in unsafe scenarios, with a higher frequency of detecting the unsafe intent from the user’s query under the current situation, and pointing out the unsafe outcomes or rejecting to answer. Meanwhile, proprietary MLLMs are not over-sensitive in safe situations; therefore, they obtain higher average safety accuracy than open-source MLLMs. Third, by comparing the performance on Chat and Embodied scenarios, we find that MLLMs all perform worse on Embodied scenarios, especially in recognizing unsafe situations. Lastly, the best-performed model, Claude 3.5 Sonnet, only scores an average accuracy of 62.2%, indicating the situation safety awareness of current MLLMs needs to be improved.
@@ 4.3 Result Diagnosis
Secondly, by comparing chat and embodied scenarios, we can find that the improvement of MLLMs on embodied tasks is very limited, even proprietary MLLMs only achieve around 58% accuracy. This shows current MLLMs have limited safety knowledge in embodied scenarios. By looking into the models’ output, we find that MLLMs often make safety judgments based on non-significant visual observations. For instance, they would judge the task of placing a knife on the table as unsafe to perform because it requires cautiousness. Also, the MLLMs sometimes can not identify the object types in the image, leading to insufficient information.
@@
Thirdly, by comparing classifying intent and query, we find that classifying the safety of intent has a significantly higher accuracy for both closed and open-source models on chat scenarios. After looking into the model’s output, we find three main error patterns due to the task of classifying the safety of the query being more complex, with the extra task of recognizing the user’s potential intent. The first one is the model ignores the unsafe situation in the image. In the example shown in Fig. 5 (middle), Gemini did not recognize the scenario is in a lab where eating might be prohibited. The second one is the model made hallucinates about safety, leading to incorrect safety judgment. For example, in Fig. 5 (left), Gemini thinks parking behind or in front of the car is dangerous without any support. The third one is the model did not follow the instructions to judge the safety of the user’s intent in the given situation. For instance, in Fig. 5 (right), llava did not judge the safety of the user’s query. Instead, it comments the user’s query in a general way.
@@ 5.1 Multi-Agent System Design
We aim to leverage our analysis results to improve the MLLM’s safety awareness when answering user’s queries. First, we introduce explicit safety reasoning, which has shown significant safety performance improvement for both chat and embodied scenarios. Second, based on our findings that more complex task settings decrease the safety judgment performance of MLLMs, we explore leveraging the multi-agent systems (Zeng et al., 2024). Specifically, we split the task of answering questions safely into several subtasks and assigned them to different MLLM agents.
@@ 5.2 Result and Analysis
We consider two baseline settings. The first one is the setting in Table. 2, where the prompt instructs the MLLMs to answer the user’s query. Second, we let MLLMs perform the intent reasoning, safety judgment, and query-responding tasks in one step. The results of our multi-agent framework are in Fig. 7, showing that the multi-agent pipeline improves the performance consistently for almost all the models in both embodied and chat subtasks. In the chat scenario, the improvements of multi-agent on the open-source models are larger, which is due to the fact they perform weaker when solving all subtasks at once. We can observe that most open-source MLLMs can not improve their performance when performing all subtasks together, compared with the base setting. With our multi-agent design, most open-source MLLMs can catch the performance of Gemini, one of the closed-source models, this shows the effectiveness of our method.
@@
Claude 65.6 76.1 83.4
GPT4o 68.7 81.7 86.0
Table 3: Investigation of MLLM’s limitation in the embodied multiagent framework by comparing performance on three settings: Setting I (Embodied Multi-Agent), Setting II (GT Environment State) and Setting III (GT Observation).
In the embodied scenario, the multi-agent design improves most of MLLMs. For open-sourced MLLMs, the improvement of CoT reasoning and multi-agent is smaller. This is because the weaker instruction-following ability makes some MLLMs unable to execute subtasks well. Also, even the performance of the best MLLM, GPT4o, is far from perfect. To investigate the reason, we perform two ablation studies on two best-performing models: GPT-4o and Claude. We replace the reasoned important environment state by the first agent with the ground truth environment state that determines the safety of a task or the ground truth observation of this environment state. The result is in Table. 3, which shows both replacements improve the performance. This means the MLLMs can not correctly locate the important environment state all the time and make visual recognition errors and hallucinations regarding safety judgment. For example, GPT-4o falsely thinks the toggling of a sink with a knife in it could cause injury and does not see the object that needs to be dropped on the floor is a cell phone. This shows that safety training in the embodied scenarios needs to be improved.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) are rapidly evolving, demonstrating impressive capabilities as multimodal assistants that interact with both humans and their environments. However, this increased sophistication introduces significant safety concerns. In this paper, we present the first evaluation and analysis of a novel safety challenge termed Multimodal Situational Safety, which explores how safety considerations vary based on the specific situation in which the user or agent is engaged. We argue that for an MLLM to respond safely—whether through language or action—it often needs to assess the safety implications of a language query within its corresponding visual context. To evaluate this capability, we develop the Multimodal Situational Safety benchmark (MSSBench) to assess the situational safety performance of current MLLMs. The dataset comprises 1,960 language query-image pairs, half of which the image context is safe, and the other half is unsafe. We also develop an evaluation framework that analyzes key safety aspects, including explicit safety reasoning, visual understanding, and, crucially, situational safety reasoning. Our findings reveal that current MLLMs struggle with this nuanced safety problem in the instruction-following setting and struggle to tackle these situational safety challenges all at once, highlighting a key area for future research. Furthermore, we develop multi-agent pipelines to coordinately solve safety challenges, which shows consistent improvement in safety over the original MLLM response. Code and data: mssbench.github.io.
@@ 1 Introduction
To comprehensively evaluate the current MLLM’s situational safety performance, we introduce a Multimodal Situational Safety benchmark (MSSBench) with 1960 language-image pairs. To assess unbalanced model behaviors, in half of the data, the image is a safe situation for answering the query, and in the other half, the image context is unsafe. Our benchmark considers two multimodal assistant scenarios: multimodal chat agents that answer users’ questions and multimodal embodied agents that plan and take actions following instructions of daily tasks. For the chat scenario, we leverage LLMs to generate candidate activities as user intents and envision an unsafe situation for these activities. Then, the examples will go through two filtering processes: LLM automatic verification and human verification performed by domain experts to ensure data quality. Finally, we prompt the LLMs to generate user queries with the intent to perform these activities. For embodied scenarios, we first manually create potentially unsafe household tasks, and define safe and unsafe situations. Then, we collect safe and unsafe visual contexts from the embodied AI simulators.
We evaluate popular open-sourced and proprietary MLLMs on the MSSBench. The results show that current MLLMs struggle with recognizing unsafe situations when answering user queries. Then, we create different evaluation variants to analyze key safety aspects of MLLMs, including explicit safety reasoning, visual understanding, and situational safety reasoning. Our main findings include: (1) Explicit safety reasoning can improve the average situational safety performance of MLLMs, but will also introduce over-sensitivity in safe situations. (2) MLLMs perform poorly in embodied scenarios due to the lack of precise visual understanding and situation safety judgment abilities. (3) Open-source MLLMs sometimes ignore crucial safety clues in the image. (4) Under settings with more subtasks, the safety performance of MLLMs decreases due to task complexity.
@@ 2 Related Work
The second one is embodied decision-making and task planning (Shridhar et al., 2020; Szot et al., 2024), which requires the MLLMs to serve as the ’brain’ of the embodied agent that plan actions for a robot to execute to complete a given household task (Driess et al., 2023; Yang et al., 2024; Li et al., 2024b; Wang et al., 2024a). This enables the MLLMs to control a robot and make it an embodied assistant. However, the improved abilities of current MLLMs on these tasks and new applications introduce new safety problems, and the safety of MLLMs under multimodal assistant scenarios has not been thoroughly studied.
@@
The generative abilities of LLMs and MLLMs carry the risk of being misused to generate harmful content. Recently, lots of efforts have been put into red-teaming MLLMs (Liu et al., 2023c; Gong et al., 2023; Shayegani et al., 2023; Qi et al., 2024; Luo et al., 2024; Shi et al., 2024). However, most of the current benchmarks study the scenarios where the language itself is clearly unsafe and leverage image modality as an attack or safety-unrelated background to trick the MLLMs into answering unsafe queries. These include using query-relevant images (Liu et al., 2023c), direct text embedding (Gong et al., 2023), and optimized adversarial images (Shayegani et al., 2023) to induce them to generate harmful responses. Moreover, Shi et al. (2024) access the safety in multimodal chat scenarios with images as additional context. Besides these, there were also concurrent efforts studying the over-sensitivity of MLLMs (Li et al., 2024c), and find that the combination of safe image and safe text inputs could be unsafe (Wang et al., 2024b). Different from existing works, we first propose a new safety problem for MLLMs in multimodal assistant applications – multimodal situational safety, where the safety of language queries varies with different visual situations. Based on this, we collect a benchmark containing chat and embodied scenarios to evaluate the MLLMs’ safety awareness. We also investigate in-depth how far we can leverage MLLMs’ capabilities to improve safety performance.
@@ 3.1 Dataset Overview
I. Physical Harm 628 32.0%
            $\bullet$ Self-harm 320 16.3%
            $\bullet$ Self-harm (Embodied Task) 120 6.0%
            $\bullet$ Other-harm 188 9.6%
II. Property Damage 876 44.7%
            $\bullet$ Public property damage 120 6.1%
            $\bullet$ Personal property damage 116 5.9%
            $\bullet$ Personal property damage (Embodied Task) 640 32.7%
III. Offensive Behavior 268 13.7%
            $\bullet$ Cultural belief violations 28 1.4%
            $\bullet$ Disruptive behaviors 148 7.3%
            $\bullet$ Religious belief infringements 92 4.7%
IV. Illegal Activities 188 9.7%
            $\bullet$ Human-restricting activities 76 3.9%
            $\bullet$ Property-restricting activities 88 4.5%
            $\bullet$ Organism-restricting activities 24 1.2%
@@ 3.3 Embodied Data Collection
We mainly consider five task categories: place an {object in hand} on a {receptacle} (Place), toggle a {receptacle} (Toggle), drop an {object in hand} (Drop), heat an {object} with microwave (Heat), pick and place {objects} (Pick&place). For each category, we define different safe and unsafe tasks by changing the objects or receptacles in the placeholder and environment states. Unsafe tasks contain uncommon environment states that could lead to unsafe outcomes. For instance, the knife in the microwave in Fig. 3. In total, we define 38 safe tasks and 38 unsafe tasks. Then, we create 5 instruction templates for each task. In total, we have $5\times(38+38)=380$ embodied instructions.
@@
After we determine the {object}, {receptacle} in the task, we run a ``$Pick\_\{object\}and\_Place\{receptacle\}"$ task defined in  Shridhar et al. (2020) with the determined {object} and {receptacle}. For the Place task and the Drop task, we randomly collect two egocentric images after the agent picks up the object and before the agent places the object. For the Toggle task, we collect an egocentric image right after the agent places the object on the receptacle from two different episodes. For Heat and Pick&place task, we collect two observations using DALL-E 3 (Betker et al., ), due to better flexibility. Therefore, we have 760 samples in total. One data example is shown in Fig. 3 (right).
@@ 3.4 Data Statistics
The Multimodal Situational Safety benchmark consists of a substantial collection of 1960 Image-Query pairs, encompassing two subsets: the embodied assistant subset, which contains 760 pairs sourced from embodied scenarios, and the chat assistant subset, comprising a larger set of 1200 pairs designed for broader situational QA scenarios. Our dataset is a balance dataset, with half of the data containing safe situations and half containing unsafe situations. The statistical details of the data in the MSSBench are presented in Table. 1.
@@ 4.1 Setup
MLLMs. The MLLMs we benchmark include both open-source models and proprietary models accessible only via API. The open-source MLLMs are: (i) LLaVA-1.6 (Liu et al., 2023b), (ii) MiniGPT4-v2 (Chen et al., 2023), (iii) Qwen-VL (Bai et al., 2023), (iv) DeepSeek (Lu et al., 2024) and (v) mPLUG-Owl2 (Ye et al., 2024). We implemented these models with their 7B version and using their default settings. For the proprietary models, we evaluated Claude 3.5 Sonnet, GPT-4o (OpenAI, 2023b), and Gemini Pro-1.5 (Reid et al., 2024).
Evaluation. For the instruction following setting, we use GPT-4o (OpenAI, 2023a) to categorize the response generated by MLLMs into safe and unsafe categories. The categories description is introduced in Tables. 4 and 5 in Sec. A.4. After categorization, we use accuracy to evaluate MLLM’s safety performance.
@@
MiniGPT-V2 98.5 2.6 50.6 98.8 0.8 49.8 48.8
Qwen-VL 96.5 3.8 50.2 99.5 0.5 50.0 50.1
mPLUG-Owl2 98.7 2.9 50.8 97.9 1.3 49.6 50.3
Llava 1.6 99.1 1.7 50.4 99.2 1.6 50.4 50.4
DeepSeek 98.6 7.8 53.2 99.7 2.4 51.1 52.4
@@ 4.2 Main Results
To begin with, we assess the performance of 8 leading multimodal large language models (MLLMs) on our MSS benchmark, the results are shown in Table. 2. To mimic the chat assistant scenario, we inform the MLLM that the image is its first-person view and the query is from a user staying with it, see the ‘Common Prompt’ in Fig. 4. The full prompt can be found in Sec. A.7. First, a common trend among all the MLLMs is that they tend to comply with and answer users’ queries in both safe and unsafe scenarios. This leads to a high safety accuracy when the situation is safe for the user’s intent and a low accuracy when the situation is unsafe. Second, comparing open-source models and proprietary models, we find that proprietary models perform better in unsafe scenarios, with a higher frequency of detecting the unsafe intent from the user’s query under the current situation, and pointing out the unsafe outcomes or rejecting to answer. Meanwhile, proprietary MLLMs are not over-sensitive in safe situations; therefore, they obtain higher average safety accuracy than open-source MLLMs. Third, by comparing the performance on Chat and Embodied scenarios, we find that MLLMs all perform worse on Embodied scenarios, especially in recognizing unsafe situations. Lastly, the best-performed model, Claude 3.5 Sonnet, only scores an average accuracy of 64.0%, indicating the situation safety awareness of current MLLMs needs to be improved.
@@ 4.3 Result Diagnosis
Secondly, by comparing chat and embodied scenarios, we can find that the improvement of MLLMs on embodied tasks is very limited, even proprietary MLLMs only achieve around 59% accuracy. We identified three main reasons from the models’ output. The first one is weaker safety knowledge in the embodied scenario, leading to more safety judgment errors. For instance, Gemini thinks dropping the phone on the floor will not cause any damage. Second, MLLMs make more visual understanding errors in the embodied scenarios and cannot identify key objects or identify them wrongly, due to MLLMs being trained less on the embodied data, and the key visual factor in the embodied scenario is more subtle. Third, MLLMs ignore key factors that may affect the safety of a task when making judgments. For example, GPT4o does not consider what object the robot is holding (a remote control) when judging whether placing it in the sink is safe.
@@
Thirdly, by comparing classifying intent and query, we find that classifying the safety of intent has a higher accuracy for both closed and open-source models. After looking into the model’s output, we see three main error patterns caused by the task of classifying the safety of the query being more complex, with the extra task of recognizing the user’s potential intent. The first one is the model ignores the unsafe situation in the image. In the example shown in Fig. 5 (middle), Gemini did not recognize the scenario is in a lab where eating might be prohibited. The second one is the model made hallucinates about safety, leading to incorrect safety judgment. For example, in Fig. 5 (left), Gemini thinks parking behind or in front of the car is dangerous without any support. The third one is the model did not follow the instructions to judge the safety of the user’s intent in the given situation. For instance, in Fig. 5 (right), llava did not judge the safety of the user’s query. Instead, it comments the user’s query in a general way.
@@ 5.1 Multi-Agent System Design
We aim to leverage our analysis results to improve the MLLM’s safety awareness when answering user’s queries. First, we introduce explicit safety reasoning, which has shown significant safety performance improvement for both chat and embodied scenarios. Second, based on our findings that more complex task settings decrease the safety judgment performance of MLLMs, we explore leveraging the multi-agent systems. Specifically, we split the task of answering questions safely into several subtasks and assigned them to different MLLM agents.
@@ 5.2 Result and Analysis
We consider two baseline settings. The first one is the setting in Table. 2, where the prompt instructs the MLLMs to answer the user’s query. Second, we let MLLMs perform the intent reasoning, safety judgment, and query-responding in one step. The results of our multi-agent framework are in Fig. 7, showing that the multi-agent pipeline improves the performance for almost all the models in both embodied and chat subtasks. In the chat scenario, the multi-agent framework significantly enhances open-source MLLMs, which struggle with solving all subtasks simultaneously due to weaker abilities. Most open-source MLLMs fail to improve in this setting, but with our multi-agent design, they can reach Gemini-level performance, demonstrating our method’s effectiveness.
@@
In the embodied scenario, the multi-agent design improves more on proprietary MLLMs. For most open-sourced MLLMs, the CoT reasoning and multi-agent do not significantly improve perfor-
@@
Claude 62.1 76.3 83.6
GPT4o 69.0 82.2 87.1
Table 3: Investigation of MLLM’s limitation in the embodied multiagent framework by comparing performance on three settings: I (Multi-Agent), II (GT Environment State), and III (GT Observation).
mance. Also, even the performance of the best MLLM, GPT4o, is far from perfect. To investigate the reason, we perform two ablation studies on two best-performing models: GPT-4o and Claude. We replace the reasoned important environment state by the first agent with the ground truth environment state that determines the safety of a task or the ground truth observation of this environment state. The result is in Table. 3, which shows both replacements improve performance. This means that MLLMs sometimes incorrectly locate the important environment state, make visual recognition errors, or have hallucinations regarding safety judgment. For example, GPT-4o falsely thinks that toggling a sink with a knife on it could cause injury and does not see that the object that needs to be dropped on the floor is a cell phone. This shows that safety training in the embodied scenarios needs to be improved.